solve 4x+x(x-2)=0

find the x intercepts of f(x)=4x+x(x-2)

Answer by robertb(5830)

4x + x(x-2) = 0
<==> x(4 + x-2) = 0
==> x(x + 2) = 0
==> x = 0, -2.